OpenStudy (anonymous):

show that integral from 0 to t f(u)du is 1/s L(s) (laplace transform

OpenStudy (turingtest):

Do mean to show that \[\int_0^\infty e^{-st}dt=\frac1s\]?

OpenStudy (anonymous):

@freddy_eighty7 Is that 1/s being mulyipling by L(s)?

OpenStudy (anonymous):

[\int\limits_{0}^{t} f(u) du = 1/s L sub f (s) \] if that makes sense

OpenStudy (turingtest):

\[\int_0^t f(u) du =\frac1s L ( f (u)) \]?

OpenStudy (anonymous):

yes